OpenStudy (anonymous):

How do i change this equation to slope intercept form 2x - 3y =-9

OpenStudy (vshiroky):

Subtract 2x from both sides... then divide both sides by -3

OpenStudy (anonymous):

simple just put the y on one side dividing by 3 2x-3y=-9 2x+9=3y on divinding by 3 2/3x+3=y

OpenStudy (anonymous):

thx